financial management

Viserion, Inc., is trying to determine its cost of debt. The firm has a debt issue outstanding with 26 years to maturity that is quoted at 95 percent of face value. The issue makes semiannual payments and has an embedded cost of 6 percent annually.
Answer #1

a) using financial calculator

n= 26*2 = 52

pv = -950

pmt = 1000*6%/2 = 30

fv =1000

compute IY = 3.20% (rounded off to two decimal place)

pretax cost of debt =3.20% x 2 =6.40%

Note- Can be solved in excel using rate function

b) after tax cost of debt =pretax cost of debt (1-tax rate)

after tax cost of debt = 6.40% (1-.25) = 4.80%

Thumbs Up Please! Thank You


answered by: ANURANJAN SARSAM